OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x – y = 3 My answer is 3 Am I right? :)

OpenStudy (uri):

4x + 2y = 12 x – y = 3 (mutiply by 2) 4x + 2y = 12 2x – 2y = 6 ========= 6x=18 x=18/6 x=3 Correct!

OpenStudy (ivettef365):

multiply the equation x-y=3 times 2 and then subtract as follows 2(x-y=3) 2x -2y = 6 now you add 4x + 2y =12 2x - 2y = 6 (now your y cancel) ----------- 6x = 18 so x = 18/6

OpenStudy (anonymous):

Thank you! :)